\begin{thebibliography}{27}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Agueh and Carlier(2011)]{agueh2011barycenters}
Martial Agueh and Guillaume Carlier.
\newblock Barycenters in the {W}asserstein space.
\newblock \emph{Society for Industrial and Applied Mathematics}, 43\penalty0
  (2):\penalty0 904--924, 2011.

\bibitem[{\'A}lvarez-Esteban et~al.(2016){\'A}lvarez-Esteban, del Barrio,
  Cuesta-Albertos, and Matr{\'a}n]{alvarez2016fixed}
Pedro~C {\'A}lvarez-Esteban, E~del Barrio, JA~Cuesta-Albertos, and
  C~Matr{\'a}n.
\newblock A fixed-point approach to barycenters in {W}asserstein space.
\newblock \emph{Journal of Mathematical Analysis and Applications},
  441\penalty0 (2):\penalty0 744--762, 2016.

\bibitem[Ambrosio et~al.(2008)Ambrosio, Gigli, and
  Savar{\'e}]{ambrosio2008gradient}
Luigi Ambrosio, Nicola Gigli, and Giuseppe Savar{\'e}.
\newblock \emph{Gradient flows: in metric spaces and in the space of
  probability measures}.
\newblock Springer Science \& Business Media, 2008.

\bibitem[Anderes et~al.(2016)Anderes, Borgwardt, and
  Miller]{anderes2016discrete}
Ethan Anderes, Steffen Borgwardt, and Jacob Miller.
\newblock Discrete {W}asserstein barycenters: Optimal transport for discrete
  data.
\newblock \emph{Mathematical Methods of Operations Research}, 84:\penalty0
  389--409, 2016.

\bibitem[Bhatia et~al.(2018)Bhatia, Jain, and Lim]{bhatia}
Rajendra Bhatia, Tanvi Jain, and Yongdo Lim.
\newblock On the {B}ures-{W}asserstein distance between positive definite
  matrices.
\newblock \emph{Expositiones Mathematicae (in press,
  https://doi.org/10.1016/j.exmath.2018.01.002)}, 2018.
\newblock ISSN 0723-0869.
\newblock \doi{https://doi.org/10.1016/j.exmath.2018.01.002}.
\newblock URL
  \url{http://www.sciencedirect.com/science/article/pii/S0723086918300021}.

\bibitem[Bonneel et~al.(2016)Bonneel, Peyr{\'e}, and
  Cuturi]{bonneel2016wasserstein}
Nicolas Bonneel, Gabriel Peyr{\'e}, and Marco Cuturi.
\newblock Wasserstein barycentric coordinates: histogram regression using
  optimal transport.
\newblock \emph{ACM Transactions on Graphics}, 35\penalty0 (4):\penalty0 71--1,
  2016.

\bibitem[Borgwardt(2017)]{borgwardt2017strongly}
Steffen Borgwardt.
\newblock Strongly polynomial 2-approximations of discrete wasserstein
  barycenters.
\newblock \emph{arXiv preprint arXiv:1704.05491}, 2017.

\bibitem[Carlier et~al.(2021)Carlier, Eichinger, and
  Kroshnin]{carlier2021entropic}
Guillaume Carlier, Katharina Eichinger, and Alexey Kroshnin.
\newblock Entropic-{W}asserstein barycenters: {PDE} characterization,
  regularity, and {CLT}.
\newblock \emph{SIAM Journal on Mathematical Analysis}, 53\penalty0
  (5):\penalty0 5880--5914, 2021.

\bibitem[Chewi et~al.(2020)Chewi, Maunu, Rigollet, and
  Stromme]{chewi2020gradient}
Sinho Chewi, Tyler Maunu, Philippe Rigollet, and Austin~J Stromme.
\newblock Gradient descent algorithms for bures-wasserstein barycenters.
\newblock In \emph{Conference on Learning Theory}, pages 1276--1304. PMLR,
  2020.

\bibitem[Cuesta-Albertos et~al.(1996)Cuesta-Albertos, Matr{\'a}n-Bea, and
  Tuero-Diaz]{cuesta1996lower}
JA~Cuesta-Albertos, C~Matr{\'a}n-Bea, and A~Tuero-Diaz.
\newblock On lower bounds for the $l_2$-{W}asserstein metric in a {H}ilbert
  space.
\newblock \emph{Journal of Theoretical Probability}, 9\penalty0 (2):\penalty0
  263--283, 1996.

\bibitem[Gramfort et~al.(2015)Gramfort, Peyr{\'e}, and
  Cuturi]{gramfort2015fast}
Alexandre Gramfort, Gabriel Peyr{\'e}, and Marco Cuturi.
\newblock Fast optimal transport averaging of neuroimaging data.
\newblock In \emph{International Conference on Information Processing in
  Medical Imaging}, pages 261--272. Springer, 2015.

\bibitem[Kroshnin et~al.(2021)Kroshnin, Spokoiny, and
  Suvorikova]{kroshnin2021statistical}
Alexey Kroshnin, Vladimir Spokoiny, and Alexandra Suvorikova.
\newblock Statistical inference for bures--wasserstein barycenters.
\newblock \emph{The Annals of Applied Probability}, 31\penalty0 (3):\penalty0
  1264--1298, 2021.

\bibitem[Masarotto et~al.(2018)Masarotto, Panaretos, and
  Zemel]{masarotto2018procrustes}
Valentina Masarotto, Victor~M Panaretos, and Yoav Zemel.
\newblock Procrustes metrics on covariance operators and optimal transportation
  of gaussian processes.
\newblock \emph{Sankhya A}, pages 1--42, 2018.

\bibitem[Masarotto et~al.(2022)Masarotto, Panaretos, and
  Zemel]{masarotto2022transportation}
Valentina Masarotto, Victor~M Panaretos, and Yoav Zemel.
\newblock Transportation-based functional anova and pca for covariance
  operators.
\newblock \emph{arXiv preprint arXiv:2212.04797}, 2022.

\bibitem[Olkin and Pukelsheim(1982)]{olkin1982distance}
Ingram Olkin and Friedrich Pukelsheim.
\newblock The distance between two random vectors with given dispersion
  matrices.
\newblock \emph{Linear Algebra and its Applications}, 48:\penalty0 257--263,
  1982.

\bibitem[Panaretos and Zemel(2020)]{panaretos2020invitation}
Victor~M Panaretos and Yoav Zemel.
\newblock \emph{An invitation to statistics in Wasserstein space}.
\newblock Springer Nature, 2020.

\bibitem[Pigoli et~al.(2014)Pigoli, Aston, Dryden, and
  Secchi]{pigoli2014distances}
Davide Pigoli, John~AD Aston, Ian~L Dryden, and Piercesare Secchi.
\newblock Distances and inference for covariance operators.
\newblock \emph{Biometrika}, 101\penalty0 (2):\penalty0 409--422, 2014.

\bibitem[Rabin et~al.(2011)Rabin, Peyr{\'e}, Delon, and
  Bernot]{rabin2011wasserstein}
Julien Rabin, Gabriel Peyr{\'e}, Julie Delon, and Marc Bernot.
\newblock Wasserstein barycenter and its application to texture mixing.
\newblock In \emph{International Conference on Scale Space and Variational
  Methods in Computer Vision}, pages 435--446. Springer, 2011.

\bibitem[Rachev and R{\"u}schendorf(1998)]{rachev1998mass}
Svetlozar~T Rachev and Ludger R{\"u}schendorf.
\newblock \emph{Mass transportation problems}.
\newblock Springer Science \& Business Media, 1998.

\bibitem[Santambrogio(2015)]{santambrogio2015optimal}
Filippo Santambrogio.
\newblock Optimal transport for applied mathematicians.
\newblock \emph{Birk{\"a}user, NY}, 55\penalty0 (58-63):\penalty0 94, 2015.

\bibitem[Santoro and Panaretos(2023)]{santoro2023large}
Leonardo~V Santoro and Victor~M Panaretos.
\newblock Large sample theory for bures-wasserstein barycentres.
\newblock \emph{arXiv preprint arXiv:2305.15592}, 2023.

\bibitem[Schmitz et~al.(2018)Schmitz, Heitz, Bonneel, Ngole, Coeurjolly,
  Cuturi, Peyr{\'e}, and Starck]{schmitz2018wasserstein}
Morgan~A Schmitz, Matthieu Heitz, Nicolas Bonneel, Fred Ngole, David
  Coeurjolly, Marco Cuturi, Gabriel Peyr{\'e}, and Jean-Luc Starck.
\newblock Wasserstein dictionary learning: Optimal transport-based unsupervised
  nonlinear dictionary learning.
\newblock \emph{SIAM Journal on Imaging Sciences}, 11\penalty0 (1):\penalty0
  643--678, 2018.

\bibitem[Seguy and Cuturi(2015)]{seguy2015principal}
Vivien Seguy and Marco Cuturi.
\newblock Principal geodesic analysis for probability measures under the
  optimal transport metric.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  3312--3320, 2015.

\bibitem[Solomon et~al.(2015)Solomon, De~Goes, Peyr{\'e}, Cuturi, Butscher,
  Nguyen, Du, and Guibas]{solomon2015convolutional}
Justin Solomon, Fernando De~Goes, Gabriel Peyr{\'e}, Marco Cuturi, Adrian
  Butscher, Andy Nguyen, Tao Du, and Leonidas Guibas.
\newblock Convolutional {W}asserstein distances: Efficient optimal
  transportation on geometric domains.
\newblock \emph{ACM Transactions on Graphics}, 34\penalty0 (4):\penalty0 1--11,
  2015.

\bibitem[Villani(2003)]{villani2003topics}
C\'{e}dric Villani.
\newblock \emph{Topics in Optimal Transportation}, volume~58.
\newblock American Mathematical Society, 2003.

\bibitem[Ye et~al.(2017)Ye, Wu, Wang, and Li]{ye2017fast}
Jianbo Ye, Panruo Wu, James~Z Wang, and Jia Li.
\newblock Fast discrete distribution clustering using {W}asserstein barycenter
  with sparse support.
\newblock \emph{IEEE Transactions on Signal Processing}, 65\penalty0
  (9):\penalty0 2317--2332, 2017.

\bibitem[Zemel and Panaretos(2019)]{zemel2019frechet}
Yoav Zemel and Victor~M Panaretos.
\newblock Fr{\'e}chet means and procrustes analysis in wasserstein space.
\newblock \emph{Bernoulli}, 25\penalty0 (2):\penalty0 932--976, 2019.

\end{thebibliography}

\documentclass[a4paper]{article}
\usepackage{hyperref}
\usepackage{url}

\usepackage{graphicx} 
\usepackage[round]{natbib}
\usepackage{algorithm}
\usepackage{algorithmic}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amsthm}
\usepackage{bm}
\usepackage{nicefrac}
\usepackage{enumerate}
\allowdisplaybreaks

\theoremstyle{plain}
\newtheorem{thm}{\protect\theoremname}
\theoremstyle{plain}

\theoremstyle{definition}

\theoremstyle{plain}

\theoremstyle{plain}

\theoremstyle{plain}

\theoremstyle{plain}
\newtheorem{lem}[]{\protect\lemmaname}
\theoremstyle{definition}

\providecommand{\algorithmname}{Algorithm}
\providecommand{\assumptionname}{Assumption}
\providecommand{\corollaryname}{Corollary}
\providecommand{\definitionname}{Definition}
\providecommand{\propositionname}{Proposition}
\providecommand{\theoremname}{Theorem}
\providecommand{\remarkname}{Remark}
\providecommand{\lemmaname}{Lemma}

\begin{document}

\title{Non-injectivity of Bures--Wasserstein barycentres in infinite dimensions}

\author{Yoav Zemel \\
Institut de Math\'ematiques \\
\'Ecole polytechnique f\'ed\'erale de Lausanne
\\
\texttt{yoav.zemel@epfl.ch}
}

\maketitle

\vspace{-0.02cm}

\begin{abstract}
We construct a counterexample to the injectivity conjecture of \cite{masarotto2018procrustes}.  Namely, we construct a class of examples of injective covariance operators on an infinite-dimensional separable Hilbert space for which the Bures--Wasserstein barycentre is highly non injective --- it has a kernel of infinite dimension.
\end{abstract}

\global\long\def\CC{\mathbb{C}}
\global\long\def\SS{S^{1}}
\global\long\def\R{\mathbb{R}}
\global\long\def\actson{\curvearrowright}
\global\long\def\ra{\rightarrow}
\global\long\def\z{\mathbf{z}}
\global\long\def\ZZ{\mathbb{Z}}
\global\long\def\NN{\mathbb{N}}
\global\long\def\sgn{\mathrm{sgn}\:}
\global\long\def\RRpos{\RR_{>0}}
\global\long\def\var{\mathrm{var}}
\global\long\def\circint{\int_{-\pi}^{\pi}}
\global\long\def\F{\mathcal{F}}
\global\long\def\pb#1{\langle#1\rangle}
\global\long\def\op{\mathrm{op}}
\global\long\def\Op{\mathrm{op}}
\global\long\def\supp{\mathrm{supp}}
\global\long\def\ceil#1{\lceil#1\rceil}
\global\long\def\TV{\mathrm{TV}}
\global\long\def\floor#1{\lfloor#1\rfloor}
\global\long\def\vt{\vartheta}
\global\long\def\vp{\varphi}
\global\long\def\class#1{[#1]}
\global\long\def\of{(\cdot)}
\global\long\def\one{\mathbbm{1}}
\global\long\def\cov{\mathrm{cov}}
\global\long\def\CC{\mathbb{C}}
\global\long\def\SS{S^{1}}
\global\long\def\RR{\mathbb{R}}
\global\long\def\actson{\curvearrowright}
\global\long\def\ra{\rightarrow}
\global\long\def\z{\mathbf{z}}
\global\long\def\ZZ{\mathbb{Z}}
\global\long\def\h{\mu}
\global\long\def\convr{*_{\RR}}
\global\long\def\x{\mathbf{x}}
\global\long\def\ve{\epsilon}
\global\long\def\cv{\mathfrak{c}}
\global\long\def\wh#1{\hat{#1}}
\global\long\def\norm#1{\left|\left|#1\right|\right|}
\global\long\def\Tmean{T}
\global\long\def\Tslope{m}
\global\long\def\degC#1{#1^{\circ}\mathrm{C}}
\global\long\def\X{\mathbf{X}}
\global\long\def\bbeta{\boldsymbol{\beta}}
\global\long\def\b{\mathbf{b}}
\global\long\def\Y{\mathbf{Y}}
\global\long\def\H{\mathbf{H}}
\global\long\def\e{\bm{\epsilon}}
\global\long\def\s{\mathbf{s}}
\global\long\def\t{\mathbf{t}}
\global\long\def\dAc{\partial A_c}
\global\long\def\cl{\mathrm{cl}}
\global\long\def\Hoe{\text{H\"older}}
\global\long\def\bb#1{\mathbb{#1}}
\global\long\def\bX{\bm{X}}
\global\long\def\bY{\bm{Y}}
\global\long\def\D{\mathfrak{D}}
\global\long\def\dGH{d_{\mathcal{GH}}}
\global\long\def\dWp{d_{\mathcal{W},p}}
\global\long\def\umu{\underline{\mu}}
\global\long\def\FLB{\mathbf{FLB}}
\global\long\def\dX{\mathbf{d_\bX}}
\global\long\def\dY{\mathbf{d_\bY}}
\global\long\def\mX{{\mu_\bX}}
\global\long\def\mY{{\mu_\bY}}	
\global\long\def\cal#1{\mathbf{\mathcal{#1}}}
\global\long\def\UXn{U_{\bX n}}
\global\long\def\dXX{{\Delta_{\bX}}}
\global\long\def\dYY{{\Delta_{\bY}}}
\global\long\def\d{\:\mathrm{d}}
\global\long\def\Ds{\mathfrak{d}}
\global\long\def\ph{\hat{p}}
\global\long\def\muh{\hat{\mu}}
\global\long\def\onevec{\one}
\global\long\def\T{\mathcal{T}}
\global\long\def\root{\mathrm{root}}
\global\long\def\parent{\mathrm{par}}
\global\long\def\children{\mathrm{children}}
\global\long\def\scp#1#2{\langle #1, #2 \rangle}
\global\long\def\X{\mathcal{X}}
\global\long\def\br{\bm{r}}
\global\long\def\bs{\bm{s}}
\global\long\def\brh{\hat{\bm{r}}}
\global\long\def\bsh{\hat{\bm{s}}}
\global\long\def\rh{\hat{r}}
\global\long\def\sh{\hat{s}}
\global\long\def\bu{{\bm{u}}}
\global\long\def\bv{\bm{v}}
\global\long\def\bh{\bm{h}}
\global\long\def\bw{\bm{w}}
\global\long\def\bD{{\bm{D}}}
\global\long\def\bx{\bm{x}}
\global\long\def\by{\bm{y}}
\global\long\def\bZ{\bm{Z}}
\newcommand{\WhnmB}{\hat{W}^{(S)}}
\global\long\def\smax{{\sigma_{\mathrm{max}}}}
\global\long\def\sT{\sigma_\T}
\global\long\def\bsT{\bm{\sigma}_\T}
\global\long\def\E{\mathcal{E}}
\global\long\def\V{\mathcal{V}}
\global\long\def\cmax{c_{\mathrm{max}}}
\global\long\def\h{h}
\global\long\def\Id{\mathrm{Id}}
\global\long\def\N{\mathcal{N}}
\global\long\def\diam{\:\mathrm{diam}}
\global\long\def\ns{{n^{*}}}
\global\long\def\ps{{p^{*}}}
\global\long\def\Wt{W^{(t)}}
\global\long\def\BL1{\mathrm{BL}_1}
\global\long\def\I{\mathcal{I}}
\global\long\def\simD{\stackrel{D}{\sim}}
\global\long\def\d{\mathrm{d}}
\global\long\def\dpp{d^p\!}
\newcommand{\innprod}[2]{{\left\langle {#1},{#2}\right\rangle}}

\section{Introduction}
Wasserstein barycenters are Fr\'echet means with respect to the Wasserstein distance of optimal transport (see e.g., the textbooks \cite{rachev1998mass}, \cite{villani2003topics},  \cite{santambrogio2015optimal}, and \cite{panaretos2020invitation}). Introduced by \cite{agueh2011barycenters}, Wasserstein barycentres have been shown to define meaningful notions of average for objects with complex geometric structures.  As such, they have found applications in numerous fields.  Examples include unsupervised dictionary learning \citep{schmitz2018wasserstein}, distributional clustering \citep{ye2017fast}, Wasserstein principal component analysis \citep{seguy2015principal}, neuroimaging \citep{gramfort2015fast} and computer vision \citep{rabin2011wasserstein,solomon2015convolutional,bonneel2016wasserstein}.

The first regularity property for Wasserstein barycenters has been already shown in \cite{agueh2011barycenters}, and states that if at least one of a finite collection of probability measures $\mu_1,\dots,\mu_n$ (with finite second moment) on $\R^d$ is absolutely continuous with a bounded density, then their barycentre $\overline \mu$ is also absolutely continuous with bounded density.  It therefore follows that the optimal transport from the barycentre $\overline\mu$ to each of the original measures $\mu_i$ is given by a transport map, even if $n-1$ of the measures possess no regularity. Perhaps surprisingly, the existence of the transport maps also holds when all the measures are finitely supported on $\R^d$ (\cite{anderes2016discrete}; see also \cite{borgwardt2017strongly}).

When the measures $\mu_1,\dots,\mu_n$ are all Gaussian, the Wasserstein distance has a closed form \citep{olkin1982distance}.  While the barycentre (known as the Bures--Wasserstein barycentre in the Gaussian case) does not admit one in the general case, it is known to be Gaussian itself \citep{agueh2011barycenters}.  Since its behaviour with respect to the mean vectors of the Gaussian distributions is trivial, it is commonplace in the Gaussian case to identify the measures with the corresponding covariances matrices (after setting all the mean vectors to be zero), so that the Bures--Wasserstein barycentre can be viewed as a barycentre of the covariance matrices.   The finite-dimensional Gaussian case of Wasserstein barycentres was already discussed by \cite{agueh2011barycenters}, and studied in more detail by a number of authors, including \cite{alvarez2016fixed} \cite{bhatia}, \cite{zemel2019frechet}, \cite{chewi2020gradient}, \cite{kroshnin2021statistical}, \cite{carlier2021entropic}, and others.

\cite{cuesta1996lower} showed that the results of \cite{olkin1982distance} are also valid when $\R^d$ is replaced by an infinite dimensional separable Hilbert space, where the analogue of covariance matrix is known as covariance operator.  \cite{masarotto2018procrustes} then studied the Bures--Wasserstein barycentre problem on separable Hilbert spaces, in the process identifying it the Procrustes distance on covariance operators considered by \cite{pigoli2014distances}.  An important question raised by \cite{masarotto2018procrustes} is whether transport maps from the Bures--Wasserstein barycentre to each of the covariance must exist, as they do in finite dimensions.  This question is not merely a mathematical curiosity; these maps are closely related to the log maps that lift measures to the tangent space \cite{ambrosio2008gradient}.  As highlighted by \cite{zemel2019frechet}, \cite{masarotto2018procrustes}, and others, this allows to apply linear techniques such as principal component analysis to the nonlinear space of Gaussian measures (or of covariance operators).  The approach employed by \cite{masarotto2018procrustes} was based on the fact, establised by \cite{cuesta1996lower}, that a transport map from $N(0,\Sigma_1)$ to $N(0,\Sigma_2)$ exists provided $\ker(\Sigma_1)\subseteq\ker(\Sigma_2)$, i.e., $\Sigma_1$ is ``more injective" than $\ker(\Sigma_2)$.  Thus \cite{masarotto2018procrustes} were led to conjecture that if all $\Sigma_1,\dots,\Sigma_n$ are all injective, then so is the barycentre $\overline \Sigma$ (see their Conjecture 17).

The injectivity conjecture is true, of course, in finite dimensions, and it also holds in infinite dimensions when the covariance operators commute ($\Sigma_i\Sigma_j=\Sigma_j\Sigma_i$).  Though they have not proven the conjecture, \cite{masarotto2022transportation} have been able to establish the existence of transport maps as bounded operators, thus guaranteeing the availability of log maps and the validity of a principal component analysis procedure.  Nevertheless, the injectivity conjecture remains open.  While no longer needed in order to employ principal component analysis, the conjecture is still relevant for large sample theory of Bures--Wasserstein barycentres; indeed, it is related to a condition recently assumed by \cite{santoro2023large} in their study of a central limit theorem for empirical Bures--Wasserstein barycenters in infinite dimensions.

This paper resloves the injectivity condition in an unequivocally negative manner.  We construct generic counterexamples that show how the conjecture is blatantly false.  The construction is, in our view, striking in its simplicity, and is all the more surprising given the existence of transport maps established by \cite{masarotto2022transportation}.  In short, we show that any sufficiently nondegenerate covariance $\Sigma$ can be the barycentre of injective covariances, even when $\Sigma$ itself is very far from being injective.  Interestingly, it is the very existence of transport maps that provides a clear path into constructing the counterexamples.

\textbf{Notation}. Below, $H$ is an infinite-dimensional separable Hilbert space, and $\Sigma$ (possibly with subscript) is a covariance operator on $H$, that is, a self-adjoint nonnegative definite trace-class linear operator from $H$ to $H$.


\section{Negative resolution of the injectivity conjecture}
\begin{thm}
Let $\Sigma$ be a covariance operator with infinite-dimensional image.  Then for $n\ge2$ there exists a collection $S_1,\dots,S_n$ of injective covariance operators for which $\Sigma$ is the (unique) barycentre.  A population version (with $S_1,\dots,S_n$ replaced by a random operator $S$) is also possible; see the end of the proof.
\end{thm}
Obviously the conclusion holds if the kernel of $\Sigma$ has a finite dimension.  This theorem is as strong as it could be, because we know that a bounded linear optimal transport map $T_i$ from $\Sigma$ to an injective operator $S_i$ exists, so that $S_i=T_i\Sigma T_i$ needs to be injective, and with $\rm{dim}(H)=\infty$ this is impossible when the range of $\Sigma$ has finite rank.

The gist of the construction is to ``create injectivity" by exhibiting a nice $T$ such $T\Sigma T$ is injective even if $\Sigma$ is not.  It is noteworthy, however, that the existence of $T$ (established in \cite{masarotto2022transportation}) is not used in our proof.
\begin{lem}
For $\Sigma$ as above, it is possible to find a bounded self-adjoint nonnegative operator $T$ such that $T\Sigma T$ is injective.
\end{lem}
\begin{proof}
It is possible to find an orthonormal basis $(\varphi_k)_{k=1}^\infty$ of $H$ such that the kernel of $\Sigma$ is contained in the span of $\varphi_1,\varphi_3,\varphi_5,\dots$.

Since $T\Sigma T$ has to be injective, $T$ must be injective, but at the same time its range must be disjoint from the kernel of $\Sigma$, the span of the odd basis functions.  It is easy to construct such an operator:  define $F:H\to H$ by $F\varphi_k=\varphi_{2k}$.  Now $F$ is not self-adjoint, but $F+F^*$ is.  This operator is not nonnegative definite, but by writing $x=\sum x_k\varphi_k$ with $\sum x_k^2=1$ and observing that
\[
\innprod {F^*x}x=
\innprod {Fx}x
=\innprod{\sum x_k\varphi_{2k}}{\sum x_i\varphi_i}
=\sum_{k=1}^\infty x_kx_{2k}
\]
is in $[-1,1]$ by the Cauchy--Schwarz inequality, we obtain that $\|F\|_\infty\le1$ and the same holds for $\|F^*\|_\infty$, where $\|\cdot\|_\infty$ denotes the operator norm.  It therefore follows that the operator\footnote{Multiplying the identity by any $c\ge2$ will do.  The relevant polynomial $1\pm ct+t^2$ has a root of magnitude $\ge1$ for any $c\in\R$, and $T$ is nonnegative definite for any $c\ge2$.}
\[
T=F+F^*+2\mathcal I
\]
is self-adjoint, nonnegative, and with $\|T\|_\infty\le 4$.

We now show that $T$ is our desired operator.  If a vector $x\in H$ satisfies $Tx\in \mathrm{ker}(\Sigma)$ then $\innprod{Tx}{\varphi_{k}}=0$ for all $k$ even.  We shall show that $x=0$.  This would imply that $\Sigma T$ is injective, and therefore $T$ is injective, so that $T\circ (\Sigma T)$ is injective, as required.

Expressing $x=\sum x_k\varphi_k$ and noting that $F^*\varphi_k=\varphi_{k/2}$ (zero if $k$ is odd), we obtain the sequence of equalities
\[
0=\innprod{Tx}{\varphi_k}
=x_{2k}+x_{k/2}+2x_k,\qquad k=2,4,6,\dots,
\]
so that
\[
x_{4k} = -2x_{2k} - x_{k},\qquad k=1,2,3,\dots.
\]
If $x\ne 0$ this implies that $x_k$ cannot go to zero, as the relevant polynomial $1+2t+t^2$ has a root of magnitude $\ge1$.  More precisely fix $k\ge1$ and let $y_j=x_{2^jk}$ so that $x_k=y_0$ and
\[
y_j = -2y_{j-1} - y_{j-2},\qquad j=2,3,\dots.
\]
Define the generating function
\[
g_k(t)=\sum_{j=0}^\infty y_jt^j = y_0 +  y_1t - \sum_{j=2}^\infty 2y_{j-1} t^{j-1}t - \sum_{j=2}^\infty y_{j-2}t^{j-2}t^2
=y_0 +y_1t - 2tg_k(t) + 2ty_0 - t^2g_k(t).
\]
Then
\[
g_k(t) = \frac{y_0+2y_0t+ y_1 t}{1+2t+t^2}
=\frac{-y_0-y_1}{(1+t)^2} + \frac {2y_0+y_1}{1+t}
= a\sum_{j=0}^\infty (j+1)(-1)^jt^j + b\sum_{j=0}^\infty (-1)^jt^j
\]
where $a=-y_0-y_1$ and $b=2y_0+y_1$ are linear functions of $y_1,y_2$.  The coefficient of $t^j$, namely $y_j$, is $(-1)^j[b + aj+a]$.  Since $y_j=x_{2^jk}\to0$ it must be that $a=0$ and $b=0$, which means that $y_0=y_1=0$ and therefore $x_k=y_0=0$.  Since $k$ is arbitrary this shows $x=0$ as desired.
\end{proof}
\begin{proof}[Proof of main theorem]
We shall use the generative model result \citep[Theorem 14]{masarotto2018procrustes} that states that if $T_1$ and $T_2$ are self-adjoint, nonnegative definite and have average identity (namely, $T_1+T_2=2\I$), then any covariance $\Sigma$ is a Bures--Wasserstein barycentre of $S_1=T_1\Sigma T_1$ and $S_2 = T_2\Sigma T_2$.

The operators
\[
T_1 = \frac 12F+ \frac12F^* + I,
\qquad T_2 = -\frac 12F - \frac 12F^* + I
\]
are bounded nonnegative self-adjoint and average to the identity so $\Sigma$ is the Fr\'echet mean of $S_1=T_1\Sigma T_1$ and $S_2=T_2\Sigma T_2$.  We have shown that $4S_1$ is injective, and injectivity of $4S_2$ can be shown in the same way, since the polynomial in the denominator of the new generating function is $(1-2t+t^2)$ which admits a root of absolute value $\ge1$.  The proof is thus complete for $n=2$.  It is easy to extend it to any finite $n\ge2$.  It is also possible to take a random variable $a$ with symmetric distribution on $[-1/2,1/2]$ and define a random bounded nonnegative self-adjoint operator $T=a(F+F^*)+\I$.  Then $\mathbb E[T]=\I$, so that $\Sigma$ is the barycentre of the random operator $T\Sigma T$.  If $\mathbb P(a=0)=0$, then $S=T\Sigma T$ is almost surely injective.
\end{proof}

\bibliographystyle{plainnat}
\bibliography{counterInj}

\end{document}